OpenStudy (anonymous):

Please help! medal You sell math textbooks and earn $6.00 an hour base pay plus 1/25 commission on your sales. You had a fantastic week working 35 hours and selling textbooks to all the math teachers in the Anoka-Hennepin school district. if your books orders totaled $7500, what would you earn a week?

OpenStudy (anonymous):

@cwrw238

OpenStudy (kittiwitti1):

List of info: -$6 per hour -35 hours -1/25 commission on your sales -TOTAL sum $7500 First, calculate how much money you would get with the $6/hr base pay.

OpenStudy (kittiwitti1):

So how much would that be?

OpenStudy (anonymous):

210$ a week right?

OpenStudy (kittiwitti1):

Yep

OpenStudy (kittiwitti1):

Sorry - I don't know how the commission factors in... I would think you take 1/25 out of $210 and then re-add it to the value... so\[\frac{210}{25}+210\]I think?

OpenStudy (kittiwitti1):

I think the $7500 is meant to throw you off

OpenStudy (kittiwitti1):

@Callisto @ganeshie8 ?

OpenStudy (anonymous):

well the answer is $510 I just can figure out how they got tht answer

OpenStudy (kittiwitti1):

hmm... let's make an equation then:\[f(x)=6.00x+\frac{6.00x}{25}, x\text{ being hours}\]and then solve for x=35:\[f(35)=6.00(35)+\frac{6.00(35)}{25}\]so how does this work out? same answer?

OpenStudy (wolf1728):

$210 is the hours times the hourly rate the commission is 1/25 of 7500 or $300

OpenStudy (kittiwitti1):

OH THAT'S WHAT A COMMISSION IS My bad D:

OpenStudy (wolf1728):

well I guess that's all settled?

OpenStudy (anonymous):

so what formula do i use?

OpenStudy (kittiwitti1):

okay so the equation is\[f(x)=6.00x+\frac{z}{25}\]x is the hours you worked and z is the money you made from selling books

OpenStudy (wolf1728):

Seems you need 2 formulas weekly pay = hours worked * hourly rate = 35 * 6.00 = $210

OpenStudy (kittiwitti1):

I combined the formulas @wolf1728

OpenStudy (kittiwitti1):

So @katlin95, plug the x and z values in, what do you get c: ?

OpenStudy (wolf1728):

commission = total sales * 1/25 commission = 7,500 * .04 = $300

OpenStudy (kittiwitti1):

or\[y=6.00(35)+\frac{7500}{25}=210+300=$510\]

OpenStudy (kittiwitti1):

Double check your answer

OpenStudy (anonymous):

wow! i see thanks!!!

OpenStudy (anonymous):

thankyou guys!!

OpenStudy (kittiwitti1):

You're welcome :) but if @wolf1728 hadn't told me what a commission was I would have gotten it wrong ^^;

OpenStudy (wolf1728):

okay katlin and thanks kitti

OpenStudy (kittiwitti1):

You're welcome :) @katlin95 do you need any more help? <:

OpenStudy (anonymous):

yes!! i have posted another one, ill tag u:)